OpenStudy (anonymous):

JUST THE ANSWER PLEASE TIME CRUNCH!!! Find the value of x so that the line passing through (-2, 7) and (x, 2) has a slope of -1/3

OpenStudy (danjs):

13

OpenStudy (misty1212):

oooh guess we gotta hurry \[\frac{2-7}{x+2}=\frac{1}{3}\]

OpenStudy (danjs):

i saw you just went though a question exactly like this, so um....

OpenStudy (danjs):

There is your answer though

OpenStudy (misty1212):

yeah sometimes it is best just to cut to the chase

OpenStudy (anonymous):

5/ x+2

OpenStudy (anonymous):

i just need more practice..

OpenStudy (misty1212):

ok \(2-7=-5\) so \[\frac{-5}{x+2}=-\frac{1}{3}\]

OpenStudy (anonymous):

this stuff makes me want to slam my head through a wall

OpenStudy (misty1212):

lol wait til you get to the hard stuff

OpenStudy (anonymous):

oh great.....

OpenStudy (misty1212):

in any case \[\frac{-5}{x+2}=-\frac{1}{3}\\ 15x=x+2\]

OpenStudy (danjs):

It is actually pretty simple if you think about what you are doing. You are just measuring how much the change in height changes vs how much the change in distance is for a line.

OpenStudy (misty1212):

lol \[15=x+2\\ x=13\] like @DanJS said

OpenStudy (danjs):

-1/3 just means, that for every 1 step down, you move over 3 steps for a line

OpenStudy (misty1212):

the problem with this or any math is that if you make a mistake at the beginning. like \(2-7=5\) you can never recover from it later it will be wrong from there on in

OpenStudy (anonymous):

ok.. kinda getting it but the more practice the better what do u guys think of this problem- Find the slope of the line that passes through (-2, 7) and (-2, 9).

OpenStudy (misty1212):

don't even try it!

OpenStudy (anonymous):

now slope is the x1, y1 thing right?

OpenStudy (misty1212):

the x coordinates are the same

OpenStudy (misty1212):

that means the line is vertical, and a vertical line does not have a slope

OpenStudy (misty1212):

so don't start to compute it ,you cannot

OpenStudy (anonymous):

@DanJS @misty1212

OpenStudy (misty1212):

you can try... you will get \[\frac{9-7}{-2+2}=\frac{2}{0}\] but that is not a number

OpenStudy (danjs):